OpenStudy (anonymous):

Find the next term of the following sequence. 25, 10, 4, ...

Nnesha (nnesha):

formula to find next term of geometric sequence \[\huge\rm r= \frac{ a_2 }{ a_1 } , \frac{ a_3 }{ a_2 } ,\frac{ a_4 }{ a_3 }\] you should get same number and for arithmetic sequence \[\huge\rm d =a_2 -a_1 \] \[a_3 - a_1 \]

OpenStudy (amoodarya):

it is geometric sequence \[a_1,a_1q,_1q^2,...\\q=\frac{a_2}{a_1}=\frac{a_3}{a_2}\\q=\frac{10}{25}=\frac{4}{10}\\q=\frac{2}{5}\\a_n=a_1q^{n-1}\\a_1=25\\a_2=10=25*\frac{2}{5}\\a_3=a_1q^2=10*\frac{2}{5}\\a_4=a_1q^3=4*\frac{2}{5}\]